# GOOD: Towards Domain Generalized Orientated Object Detection

### v1

@@ 3.2 CLIP-driven Style Hallucination
When training on the source domain, a domain generalized oriented object detector is supposed to encounter as many styles as possible. CLIP [36] learns image representation from webwide image-text pairs, which have seen extensive styles in the training stage. The abundant style information could benefit an oriented object detector when generalized to unseen target domains that have unseen styles. To this end, we propose a novel CLIP-driven style hallucination scheme to enrich the style diversity when to learn a generalized oriented object detector.
As an object detector usually leverages features from each stage of a backbone to represent the region of interest, it is necessary to augment the style variation on the features from each stage. Specifically, let $f_{\mathcal{I}}$, $f_{\mathcal{T}}$ and $f$ denote the CLIP image encoder, CLIP text encoder and the backbone of an object detector, respectively. From shallower to deeper, assume $\mathbf{F}_{\mathcal{I}}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ and $\mathbf{F}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ denote the feature representation from the $i^{th}$ block ($i=1,\cdots,4$) of the CLIP image encoder and the detector backbone, respectively, given by
@@
Take the feature representation $\mathbf{F}_{\mathcal{I}}^{i}$ from CLIP image encoder as an example. Following existing style-transfer methods [20, 7], the channel-wise normalization mean $\boldsymbol{\mu}_{\mathcal{I}}^{i}$ and standard deviation $\boldsymbol{\sigma}_{\mathcal{I}}^{i}$ carries the style of an image $\mathbf{I}$, computed as
@@
Similarly, let $\boldsymbol{\mu}^{i}$ and $\boldsymbol{\sigma}^{i}$ denote the mean and standard deviation of the feature representation $\mathbf{F}^{i}$ from the detector backbone $f$. Then, the style-diversified counterpart $\mathbf{\widetilde{F}}^{i}$ from $\mathbf{F}^{i}$, which carries the style information from CLIP encoder, can be readily used by AdaIN [20], given by
@@
After that, both the feature representation from the backbone $\mathbf{F}^{i}$ and after style hallucination $\mathbf{\widetilde{F}}^{i}$ ($i=1,\cdots,4$) are processed by a feature pyramid network (FPN) to generate the fused feature representation before and after style hallucination, denoted as $\mathbf{F}$ and $\mathbf{\widetilde{F}}$, respectively.
@@ 3.3 Rotation-aware Consistency Learning
Even if plenty of hallucinated styles can be assumed, learning a domain generalized oriented object detector is still particularly challenging, as the orientation information of the content under varied styles need to keep consistent. Thus, we propose a rotation-aware consistency (RAC) learning to handle this issue. As the activation patterns inevitable shift on the target domains and yield less discriminative RoIs (discussed in Sec. 1), we consider to pose RAC between the RoIs between the original samples and the style-hallucinated samples.
Given a typical two-stage oriented object detector, after extracting feature representation $\mathbf{F}^{4}$ from the backbone, the regional proposal network (RPN) generates $n$ horizontal region of interests (HRoIs) $\{\mathcal{H}_{i}\}$, where $i=1,\cdots,n$. Generally, $\mathcal{H}_{i}$ is under the format of $\{\overline{x},\overline{y},\overline{w},\overline{h}\}$, which describe the 2D location of a RoI. Then, $\{\mathcal{H}_{i}\}$ need to be mapped to rotated RoIs (RRoIs) $\{\mathcal{R}_{i}\}$ to represent the orientation information, where $\mathcal{R}_{i}$ is under the format of $(x,y,w,h,\theta)$.
@@
The HRoIs’ feature representation of the original image $\mathbf{F}_{\mathcal{H}}$ and the style hallucinated image $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are learnt from $\mathbf{F}$ and $\mathbf{\widetilde{F}}$ by region proposal and feature alignment, respectively. To maintain the content consistency before and after style hallucination, both HRoIs’ representations $\mathbf{F}_{\mathcal{H}}$ and $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are projected into the embeddings $\mathbf{z}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ by two MLPs $f_{1}(\cdot)$ and $f_{2}(\cdot)$, respectively. Here $d$ refers to the dimension size of each embedding.
@@
where $\sigma(\cdot)$ is a binary function determined by the mean intersection over union (mIoU) value between a region of interest (e.g., $\mathcal{H}_{j}$) and ground truth (denoted as $\rm{gt}$), given by
@@
To realize this objective, two MLPs $f_{3}(\cdot)$ and $f_{4}(\cdot)$ are used to project $\mathbf{F}_{\mathcal{R}}$ and $\mathbf{\widetilde{F}}_{\mathcal{R}}$ into embeddings, which we denote as $\mathbf{z}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$. Similarly, we have $\mathbf{z}^{\prime}=[\mathbf{z}_{\mathcal{R}},\mathbf{\widetilde{z}}_{\mathcal{R}}]\in\mathbb{R}^{d\times 2n}$, and let $z^{\prime}_{j},z^{\prime}_{k}\in\mathbf{z}^{\prime}$, where $z^{\prime}_{j},z^{\prime}_{k}\in\mathbb{R}^{d\times 1}$, $j,k=1,\cdots,2n$ and $j\neq k$. Assume $z^{\prime}_{j,+}$ and $z^{\prime}_{j}$ construct a positive pair, and the rotation-aware consistency learning objective for RRoIs (denoted as $\mathcal{L}_{RAC}$) can be computed as
@@ 3.4 Style Consistency Learning
For oriented objects, a domain generalized detector is supposed to show robust orientation prediction despite the great style variation. During its training, it is necessary to constraint the style hallucination to be non-geometry, so that the original samples and the style-hallucinated samples could carry the same structural information. We propose a style consistency (SEC) learning to realize this objective.
@@
Both category-level scores $\mathbf{p}$ and $\mathbf{\widetilde{p}}$ need to be restricted in a way such that the pixel-level category remain unchanged in both original and style diversified samples. The Jensen-Shannon Divergence (JSD) is utilized to realize this style consistency objective, given by
@@ 3.5 Implementation Details
$\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{reg}+\mathcal{L}_{HC}+\mathcal{L}_{RAC}+\mathcal{L}_{SEC}.$ (11)
@@ IV DOTA-V1 SODA, FAIR1M, HRSC 1 ship
Table 1: Experimental settings for domain generalized oriented object detection. Following prior generalized driving-scene detection works [46, 29, 42], the shared object categories among different datasets are used to evaluate the domain generalized oriented object detector.
@@
RetinaNet OBB [30] 28.78 28.31 26.34 37.82 37.72 35.08 39.39 29.65
FCOS OBB [41] 30.47 31.46 29.76 37.49 39.80 37.45 42.46 31.86
ATSS OBB [59] 35.89 31.12 29.39 37.38 38.30 36.41 41.12 34.67
R3Det [51] 35.10 34.79 33.79 38.77 39.33 37.39 44.11 36.51
SA2Net [15] 32.31 32.57 31.41 35.21 35.70 33.37 33.25 29.90
GWD [52] 38.59 39.56 37.35 38.70 39.77 38.62 46.46 40.31
KLD [53] 37.59 36.76 34.92 39.19 41.19 39.14 48.57 36.99
SASM [18] 35.69 35.83 33.82 41.20 42.28 40.15 47.34 38.15
RepPoint [26] 37.80 36.86 34.85 37.41 40.53 37.66 50.49 40.63
@@
FR OBB [37] 33.15 33.17 30.79 34.98 36.89 34.89 44.14 34.46
MR OBB [17] 34.46 35.99 33.30 39.81 42.60 41.08 46.32 36.54
Oriented RCNN [48] 35.72 34.97 34.19 40.07 41.23 38.35 48.83 37.34
RoI-Tran [11] 36.82 37.02 35.21 41.93 43.19 40.92 52.06 41.18
ReDet [16] 40.46 41.06 38.35 42.18 44.25 41.05 54.85 43.09
@@
Table 2: Result comparison between the proposed GOOD and other oriented object detectors on unseen target domains. Evaluation metric $mAP_{50}$ (in %). For fair evaluation, all the compared methods use the ResNet-50 backbone pre-trained on ImageNet.
@@ 4.1 Dataset & Evaluation Protocols
FAIR1M [39] is a fine-grained aerial object detection dataset with 37 corresponding fine-grained categories. It has more than 15,000 images and 1 million instances. To set up domain generalization setting which exists same object categories from cross-domain, we merge the 11 fine-grained airplane categories into a single airplane category, 9 fine-grained ship categories into a ship category, 10 fine-grained vehicle categories into both large vehicle and small vehicle categories. On the other hand, as the test set of FAIR1M is not publicly available, we randomly split the official training set as 8:2 as the source domain for training and target domain for inference, respectively.
@@
DOTA [47] consists of 2,806 aerial images with 188,282 instances. The annotation of DOTA  [^2] https://captain-whu.github.io/DOTA/index.html has three versions, namely, DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 [12]. DOTA-V1.0 has 15 object categories. DOTA-V1.5 and DOTA-V2.0 has one and three more object categories than DOTA-V1.0, respectively. Following the default setting of MMRotate  [^3] https://github.com/open-mmlab/mmrotate, the images are split into 1024$\times$1024 for training, and are merged for evaluation.
@@
SODA [8] consists of 2,513 images with up to 872,069 instances from 9 classes. It is an oriented object detection benchmark for small objects. All the data split and pre-processing follow the default setting of MMRotate  [^4] https://github.com/shaunyuan22/SODA-mmrotate.
@@
HRSC [31] is a ship detection dataset with arbitrary orientations. The training and testing split directly follows the default setting of MMRotate, where all the images are resized to 800$\times$800 for both training and testing.
Evaluation Protocols.
@@
Following prior domain generalized driving-scene object detection protocols [46, 29, 42], the shared object categories among different datasets are used to evaluate the domain generalized object detector. The four domain generalization settings and the corresponding object categories are listed in Table 1.
@@
Evaluation Metric. For all the experiments, Mean Average Precision (mAP) is used as the evaluation metric. In line with [47], mAP@0.5 is reported, which classifies a prediction as a true positive when it aligns with the ground-truth label and achieves an Intersection over Union (IOU) score of over 0.5 with the ground-truth bounding box.

### v2

@@ 3.2 CLIP-driven Style Hallucination
When training on the source domain, a domain generalized oriented object detector is supposed to encounter as many styles as possible. CLIP Radford et al. (2021) learns image representation from webwide image-text pairs, which have seen extensive styles in the training stage. The abundant style information could benefit an oriented object detector when generalized to unseen target domains that have unseen styles. To this end, we propose a novel CLIP-driven style hallucination scheme to enrich the style diversity to learn a generalized oriented object detector.
As an object detector usually leverages features from each stage of a backbone to represent the region of interest, it is necessary to augment the style variation on the features from each stage. Specifically, let $f_{\mathcal{I}}$ and $f$ denote the CLIP image encoder and the backbone of an object detector, respectively. The CLIP image encoder is ResNet-50. From shallower to deeper, assume $\mathbf{F}_{\mathcal{I}}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ and $\mathbf{F}^{i}\in\mathbb{R}^{W_{i}\times H_{i}\times C_{i}}$ denote the feature representation from the $i^{th}$ block ($i=1,\cdots,4$) of the CLIP image encoder and the detector backbone, respectively, given by
@@
Take the feature representation $\mathbf{F}_{\mathcal{I}}^{i}$ from CLIP image encoder as an example. Following existing style-transfer methods Huang and Belongie (2017); Chen et al. (2021), the channel-wise normalization mean $\boldsymbol{\mu}_{\mathcal{I}}^{i}$ and standard deviation $\boldsymbol{\sigma}_{\mathcal{I}}^{i}$ carries the style of an image $\mathbf{I}$, computed as
@@
Similarly, let $\boldsymbol{\mu}^{i}$ and $\boldsymbol{\sigma}^{i}$ denote the mean and standard deviation of the feature representation $\mathbf{F}^{i}$ from the detector backbone $f$. Then, the style-diversified counterpart $\mathbf{\widetilde{F}}^{i}$ from $\mathbf{F}^{i}$, which carries the style information from CLIP encoder, can be readily used by AdaIN Huang and Belongie (2017), given by
@@
After that, both the feature representation from the backbone $\mathbf{F}^{i}$ and after style hallucination $\mathbf{\widetilde{F}}^{i}$ ($i=1,\cdots,4$) are processed by a feature pyramid network (FPN) to learn the fused feature representation before and after style hallucination, which are denoted as $\mathbf{F}$ and $\mathbf{\widetilde{F}}$, respectively.
@@ 3.3 Rotation-aware Consistency Learning
Even if the hallucinated style can be assumed, learning a domain generalized oriented object detector is still particularly challenging, as the orientation information of the content under varied styles need to keep consistent. Thus, we propose a rotation-aware consistency (RAC) learning to handle this issue. As the activation patterns inevitably shift on the target domains and yield less discriminative RoIs (discussed in Sec. 1), we consider to pose RAC between the RoIs from the original samples and the RoIs from the style-hallucinated samples.
Given a typical two-stage oriented object detector, after extracting feature representation $\mathbf{F}^{4}$ from the backbone, the regional proposal network (RPN) generates $n$ horizontal region of interests (HRoIs) $\{\mathcal{H}_{i}\}$, where $i=1,\cdots,n$. Generally, $\mathcal{H}_{i}$ is under the format of $\{\overline{x},\overline{y},\overline{w},\overline{h}\}$, which describes the 2D location of a RoI. Then, $\{\mathcal{H}_{i}\}$ need to be mapped to rotated RoIs (RRoIs) $\{\mathcal{R}_{i}\}$ to represent the orientation information, where $\mathcal{R}_{i}$ is under the format of $(x,y,w,h,\theta)$.
@@
The HRoIs’ feature representation of the original image $\mathbf{F}_{\mathcal{H}}$ and the style hallucinated image $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are learned from $\mathbf{F}$ and $\mathbf{\widetilde{F}}$ by region proposal and feature alignment, respectively. To maintain the content consistency before and after style hallucination, both HRoIs’ representations $\mathbf{F}_{\mathcal{H}}$ and $\mathbf{\widetilde{F}}_{\mathcal{H}}$ are projected into the embeddings $\mathbf{z}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{H}}\in\mathbb{R}^{d\times n}$ by two MLPs $f_{1}(\cdot)$ and $f_{2}(\cdot)$, respectively. Here $d$ refers to the dimension size of each embedding.
@@
where $\tau$ is a temperature hyper-parameter, and is by default set to be 1.0. $\sigma(\cdot)$ is a binary function determined by the mean intersection over union (mIoU) value between a region of interest (e.g., $\mathcal{H}_{j}$) and ground truth (denoted as $\rm{gt}$), given by
@@
To realize this objective, two MLPs $f_{3}(\cdot)$ and $f_{4}(\cdot)$ are used to project $\mathbf{F}_{\mathcal{R}}$ and $\mathbf{\widetilde{F}}_{\mathcal{R}}$ into embeddings, which we denote as $\mathbf{z}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$ and $\mathbf{\widetilde{z}}_{\mathcal{R}}\in\mathbb{R}^{d\times n}$. Similarly, we have $\mathbf{z}^{\prime}=[\mathbf{z}_{\mathcal{R}},\mathbf{\widetilde{z}}_{\mathcal{R}}]\in\mathbb{R}^{d\times 2n}$, where $z^{\prime}_{j},z^{\prime}_{k}\in\mathbf{z}^{\prime}$, $z^{\prime}_{j},z^{\prime}_{k}\in\mathbb{R}^{d\times 1}$, $j,k=1,\cdots,2n$ and $j\neq k$. Assume $z^{\prime}_{j,+}$ and $z^{\prime}_{j}$ construct a positive pair, and the rotation-aware consistency learning objective for RRoIs (denoted as $\mathcal{L}_{RAC}$) can be computed as
@@
While $\mathcal{L}_{HCL}$ in Eq. 6 is also applicable to predict horizontal bounding boxes on unseen target domains, the proposed $\mathcal{L}_{RAC}$ is especially devised for oriented bounding boxes and is able to constrain the orientation information of an oriented object detection. Notably, it allows the pre- and post- hallucinated HRoIs and RRoIs that have the most similar orientation information to be aligned, while the rest are pushed as apart as possible.
@@ 3.4 Style Consistency Learning
For oriented objects, a domain generalized detector is supposed to show robust orientation prediction despite the great style variation. During its training, it is necessary to constrain the style hallucination to be non-geometry, so that the original samples and the style-hallucinated samples could carry the same structural information. We propose a style consistency (SEC) learning to realize this objective.
@@
Both category-level scores $\mathbf{p}$ and $\mathbf{\widetilde{p}}$ need to be restricted in a way such that the pixel-level category remains unchanged in both original and style diversified samples. For conventional restrictions such as $l$-2 norm and Kullback-Leibler (KL) divergence, the detector training can become unstable when the numerical value between $\mathbf{p}$ and $\mathbf{\widetilde{p}}$ varies greatly after hallucination. To address this issue, the Jensen-Shannon Divergence (JSD) Menéndez et al. (1997) is utilized to realize this style consistency objective, given by
@@ 3.5 Implementation Details
$\mathcal{L}=\mathcal{L}_{cls}+\mathcal{L}_{reg}+\mathcal{L}_{HCL}+\mathcal{L}_{RAC}+\mathcal{L}_{SEC}.$ (11)
@@
Table 1: Experimental settings for domain generalized oriented object detection. Following prior generalized driving-scene detection works Wu and Deng (2022); Lin et al. (2021); Vidit et al. (2023), the shared object categories among different datasets are used to evaluate the domain generalized oriented object detector.
@@ IV DOTA-V1 SODA, FAIR1M, HRSC 1 ship
Table 2: Result comparison between the proposed GOOD and existing oriented object detectors on unseen target domains and cross-domain object detectors. Evaluation metric $mAP_{50}$ (in %). For fair evaluation, all the compared methods use the ResNet-50 backbone pre-trained on ImageNet. Top three results are highlighted as best, second and third, respectively.
@@
RetinaNet OBB Lin et al. (2017) 28.78 28.31 26.34 37.82 37.72 35.08 39.39 29.65
FCOS OBB Tian et al. (2020) 30.47 31.46 29.76 37.49 39.80 37.45 42.46 31.86
ATSS OBB Zhang et al. (2020) 35.89 31.12 29.39 37.38 38.30 36.41 41.12 34.67
R3Det Yang et al. (2021b) 35.10 34.79 33.79 38.77 39.33 37.39 44.11 36.51
SA2Net Han et al. (2021a) 32.31 32.57 31.41 35.21 35.70 33.37 33.25 29.90
GWD Yang et al. (2021c) 38.59 39.56 37.35 38.70 39.77 38.62 46.46 40.31
KLD Yang et al. (2021d) 37.59 36.76 34.92 39.19 41.19 39.14 48.57 36.99
SASM Hou et al. (2022) 35.69 35.83 33.82 41.20 42.28 40.15 47.34 38.15
RepPoint Li et al. (2022c) 37.80 36.86 34.85 37.41 40.53 37.66 50.49 40.63
PSCD Yu and Da (2023) 33.59 32.41 30.86 40.42 39.65 36.70 43.38 34.72
DCFL Xu et al. (2023) 39.85 39.07 38.15 41.69 40.56 38.72 47.88 39.63
ACM Xu et al. (2024) 36.17 36.58 33.74 38.06 38.47 35.89 45.09 36.84
@@
FR OBB Ren et al. (2015) 33.15 33.17 30.79 34.98 36.89 34.89 44.14 34.46
MR OBB He et al. (2017) 34.46 35.99 33.30 39.81 42.60 41.08 46.32 36.54
Oriented RCNN Xie et al. (2021) 35.72 34.97 34.19 40.07 41.23 38.35 48.83 37.34
RoI-Tran Ding et al. (2019) 36.82 37.02 35.21 41.93 43.19 40.92 52.06 41.18
ReDet Han et al. (2021b) 40.46 41.06 38.35 42.18 44.25 41.05 54.85 43.09
ARC Pu et al. (2023) 34.86 35.07 33.42 36.51 37.96 35.74 45.25 34.78
JDet Xiao et al. (2024) 37.05 37.28 35.87 38.46 39.07 37.21 47.63 38.97
Cross-Domain Detector:
S-DGOD Wu and Deng (2022) 29.37 30.02 27.86 36.80 36.51 34.96 38.05 30.67
DomainGen Vidit et al. (2023) 35.49 35.08 34.12 42.05 41.86 40.57 43.41 37.26
SFOD-RS Liu et al. (2024a) 34.56 34.19 33.74 40.47 39.63 38.85 41.73 33.40
PDGA Liu et al. (2024b) 36.05 37.13 35.22 41.97 42.58 39.24 40.52 32.85
@@ 4.1.1 Datasets
FAIR1M Sun et al. (2022) is a fine-grained aerial object detection dataset with 37 corresponding fine-grained categories. It has more than 15,000 images and 1 million instances. To set up a domain generalization setting which exists same object categories from cross-domain, we merge the 11 fine-grained airplane categories into a single airplane category, 9 fine-grained ship categories into a ship category, 10 fine-grained vehicle categories into large vehicle and small vehicle categories, respectively. On the other hand, as the test set of FAIR1M is not publicly available, we randomly split the official training set into an 8:2 ratio, designating the source domain for training and the target domain for inference, respectively. The images are split into 1024$\times$1024 for training, and are merged for evaluation. The overlap between two cropped image patches is 200$\times$200.
@@
DOTA Xia et al. (2018) consists of 2,806 aerial images with 188,282 instances. The annotation of DOTA  [^2] https://captain-whu.github.io/DOTA/index.html has three versions, namely, DOTA-V1.0, DOTA-V1.5 and DOTA-V2.0 Ding et al. (2021). DOTA-V1.0 has 15 object categories. DOTA-V1.5 and DOTA-V2.0 has one and three more object categories than DOTA-V1.0, respectively. Following the default setting of MMRotate  [^3] https://github.com/open-mmlab/mmrotate, the images are split into 1024$\times$1024 for training with an overlap of 200$\times$200, and are merged for evaluation.
@@
SODA Cheng et al. (2023) consists of 2,513 images with up to 872,069 instances from 9 classes. It is an oriented object detection benchmark for small objects. All the data split and pre-processing follow the default setting of MMRotate  [^4] https://github.com/shaunyuan22/SODA-mmrotate. The patch resolution is 1024$\times$1024, and the overlap is 200$\times$200.
@@
HRSC Liu et al. (2016) is a ship detection dataset with arbitrary orientations. The training and testing split directly follows the default setting of MMRotate, where all the images are resized to 1024$\times$1024 for both training and testing.
4.1.2 Evaluation Protocols
@@
The object category definition and number between SODA Cheng et al. (2023), FAIR1M Sun et al. (2022), DOTA Xia et al. (2018) and HRSC Liu et al. (2016) is different. Some categories in FAIR1M Sun et al. (2022) do not occur in the rest three datasets, and vice versa. Following prior domain generalized driving-scene object detection Wu and Deng (2022); Lin et al. (2021); Vidit et al. (2023), the common object categories among different datasets are used to evaluate the domain generalized object detector. The four generalization settings and the corresponding object categories are listed in Table 1.
Specifically, the object category number and the object categorization system is different among these four domains. The common object category number is 10, 7, 4 and 1, when we involve two, two, three and four of the oriented object detection datasets, namely, DOTA, SODA, FAIR1M and HRSC. To highlight, if all four domains are involved, there is only one common category, which is too specialized for ship detection. If three domains are involved, there are only four common categories. They are still much less than the in-domain object detection tasks, where there are usually tens of object categories. Therefore, a trade-off between domain diversity and object category number has to be made. The first and second settings do a trade-off for more common object categories from three datasets, while the third and fourth settings do a trade-off for more domain diversity, including all four datasets.
@@
For all the experiments, Mean Average Precision (mAP) is used as the evaluation metric. In line with Xia et al. (2018), mAP@0.5 is reported, which classifies a prediction as a true positive when it aligns with the ground-truth label and achieves an Intersection over Union (IOU) score of over 0.5 with the ground-truth bounding box.